OpenStudy (anonymous):

I'm a newbie here! Please help me. Find two real numbers whose sum is 24 and whose product is a minimum.

terenzreignz (terenzreignz):

Sure ^_^ Let's let those two numbers be x and y for now... ok?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Given: x + y = 24 Minimize: xy Or xy = x(24 - x)...substituting 24 - x for y. So we want to minimize x(24 - x) = 24x - x^2 Take its derivative (to find minimum), you get 24 - 2x Set equal to zero to find critical points... 24 - 2x = 0 x = 12...is a critical point. However x = 12 is a rel max, so x = 12 and hence y = 12 will be the maximum values of x and y that will maximize xy or a product of 144. This function has no minimum.

OpenStudy (anonymous):

Thank you so much for helping me! :) @Easyaspi314

OpenStudy (anonymous):

welcome. I would think that you would be asked to find values of x and y that will maximize xy, given that x+y = 24.

OpenStudy (anonymous):

OK :) @Easyaspi314